Ben is twice as old as james, who in turn is 5 years older than Elvin. If their total age is 47 years how old is each of them.

Found 2 solutions by josgarithmetic, saw:
Answer by josgarithmetic(39617)

Ben          2v+10
James         v+5
Elvin          v
SUM           47

Very simple algebra, Elvin is 8.
.
.

Answer by saw(34)

Solution.
Let James's age be x, Ben's be 2x, and Elvin's be x-5
From the condition we have this equation
x+2x+(x-5) = 47
Remove parenthesis
x+2x+x-5 = 47
Collect the like terms
4x-5 = 47
Move constant to the right
4x = 47+5
Add the numbers
4x = 52
Divide both sides by 4
x = 13
So James is 13 years old.
Ben is 2*13 = 26 years old.
Elvin is 13-5 = 8 years old.
Check-----
x+2x+(x-5) = 47
13+26+8 = 47
47 = 47